One person can do a certain job in one-half hour, and another person can do the same job in one hour and ten minutes. How many minutes will they take to do the job together?
Answer by nerdybill(7384)

One person can do a certain job in one-half hour, and another person can do the same job in one hour and ten minutes. How many minutes will they take to do the job together?
one-half hour = .5 hours
one hour and ten minutes = 1+10/60 = 1+1/6 = 7/6 hours
.
Let t = time (hours) it takes for both
then
t(1/.5 + 1/(7/6)) = 1
t(1/.5 + 6/7) = 1
t(1/.5 + 6/7) = 1
multiply both sides by 3.5
t(7 + 3) = 3.5
t(10) = 3.5
t = .35 hour
or
t = .35 *60 = 21 minutes